# Exploring the Effects of User-Agent and User-Designer Similarity in Virtual Human Design to Promote Mental Health Intentions for College Students

### v1

@@ 6. Limitations
In addition to the limitations posed by the platform’s catalog, the first part of our user study also faced challenges with the diversity and balance of its sample population. Although the number of participants was increased from previous efforts (Feijóo-García et al., 2023) and had a larger population of female participants (n=182), the group remained predominantly male (n=281) and also lacked sufficient representation from some ethnic backgrounds. Our user study depended on the student population we could recruit from our institution at the time and was limited to its socio-geographic context. We acknowledge that we could have recruited a more balanced demographic sample through platforms that allow demographic recruitment specifications (e.g., Prolific (Palan and Schitter, 2018)). However, we were limited by our resources and relied on a sample population (n=481) originating from the University of Florida. Also, we acknowledge that the decision to include only computer science students could limit the broader applicability of the findings, as it does not represent a cross-section of the general population. Please keep in mind that we limited our population to this context as previous literature suggests computer science students experience the highest rates of anxiety and depression among STEM college students (Danowitz and Beddoes, 2018), with female computer science students being particularly more at risk (Rosenstein et al., 2020; Soares Passos et al., 2020). We recommend considering these limitations to interpret our findings and observations better.
@@
Another possible limitation concerns the reuse of the same sample of participants across different stages of the study, which could have introduced biases, especially in the second part of the user study: Their previous experiences designing the virtual agents might have influenced their subsequent perceptions of the agents and their designers. Moreover, although we gave at least four days between the parts of the study, this period might not have been sufficient to prevent carryover effects, potentially affecting the outcomes. These limitations should be considered when interpreting our observations, and they may help explore further research in this area.
@@
The study also did not sufficiently explore the influence of factors beyond gender and ethnicity, such as facial expressions and physical attractiveness, on participants’ perceptions of virtual humans. For example, the top-selected male virtual humans predominantly displayed happy expressions, which might have influenced participant choices and highlighted a potential oversight regarding the impact of facial cues. When interpreting our observations, we should acknowledge these limitations and consider their implications for the study’s design and outcomes. By doing so, future research can be better positioned to address these shortcomings, which may help explore further research in this area.
@@
Moreover, in the second part of the study, it was observed that participants from the Intervention group - “Explicit legacy” (Section 4.2.1) struggled to fully recall the designer’s demographic characteristics. This suggests that future research should explore alternative methods to mitigate this issue. However, this limitation also provided valuable insights: designer cues should not be limited to verbal messages. These findings showed that the visual demographics of the agent could overshadow the verbally communicated demographics of the designer, indicating that the design process should carefully consider the balance between these two forms of communication.
These limitations provide helpful insights for future research to address the complexity involved in designing virtual humans and understanding the impact of their demographic characteristics on user perceptions.
@@ 7. Conclusions
Finally, virtual humans were not found to be more effective in promoting mental health conversations when co-designed by computer science students with similar demographic characteristics to their end users. However, observations indicate that the visual cues of the virtual humans overshadowed the information verbally communicated about the designer’s demographics, possibly affecting participants’ intentions toward mental health conversations. Thus, the demographic information of the designers was not sufficient to influence participants’ intentions toward mental health conversations, suggesting that tailoring designer cues in virtual human design is a complex task that should be explored further, especially in sensitive scenarios such as mental health support.

### v2

@@ 6. Limitations
In addition to the limitations posed by the platform’s catalog, the first part of our user study also faced challenges with the diversity and balance of its sample population. Although the number of participants was increased from previous efforts (Feijóo-García et al., 2023) and had a larger population of female participants (n=182), the group remained predominantly male (n=281) and also lacked sufficient representation from some ethnic backgrounds. We acknowledge that we could have recruited a more balanced demographic sample through platforms that allow demographic recruitment specifications (e.g., Prolific (Palan and Schitter, 2018)). However, our user study depended on the student population we could recruit from our institution at the time and was limited to its socio-geographic context: we were limited by our resources and relied on a sample population (n=481) originating from the University of Florida. Also, we acknowledge that the decision to include only computer science students could limit the broader applicability of the findings, as it does not represent a cross-section of the general population. Please keep in mind that we limited our population to this context as previous literature suggests computer science students experience the highest rates of anxiety and depression among STEM college students (Danowitz and Beddoes, 2018), with female computer science students being particularly more at risk (Rosenstein et al., 2020; Soares Passos et al., 2020).
@@
Another possible limitation concerns the reuse of the same sample of participants across different stages of the study, which could have introduced biases, especially in the second part of the user study: Their previous experiences designing the virtual agents might have influenced their subsequent perceptions of the agents and their designers. Moreover, although we gave at least four days between the parts of the study, this period might not have been sufficient to prevent carryover effects, potentially affecting the outcomes.
@@
The study also did not sufficiently explore the influence of factors beyond gender and ethnicity, such as facial expressions and physical attractiveness, on participants’ perceptions of virtual humans. For example, the top-selected male virtual humans predominantly displayed happy expressions, which might have influenced participant choices and highlighted a potential oversight regarding the impact of facial cues.
@@
Moreover, in the second part of the study, it was observed that participants from the Intervention group - “Explicit legacy” (Section 4.2.1) struggled to fully recall the designer’s demographic characteristics, possibly affecting participants’ intentions toward mental health conversations. This suggests that future research should explore alternative methods to mitigate this issue and consider that information on designer cues should not be limited to verbal messages.
@@ 7. Conclusions
Finally, virtual humans were not found to be more effective in promoting mental health conversations when co-designed by computer science students with similar demographic characteristics to their end users. However, as mentioned above, the visual cues of the virtual humans overshadowed the information verbally communicated about the designer’s demographics, possibly affecting participants’ intentions toward mental health conversations. Thus, the designers’ demographic information was insufficient to influence participants’ intentions toward mental health conversations. This suggests that tailoring designer cues in virtual human design is a complex task that should be explored further, especially in sensitive scenarios such as mental health support.